In a vertical frictionless piston and cylinder arrangement, 10 grams of water are 

heated to expand from 5 bar and 200oC to 500oC by a heat reservoir at 600oC. 

Find work and heat exchanged by water during this process. Find entropy 

change of water, heat reservoir and universe.


1
Expert's answer
2021-07-08T13:26:36-0400

Gives

"M_1=10gm=0.01kg"


"P_1=5 bar,T_1=200\u00b0c,T_2=500\u00b0c,T'=600\u00b0c"



"Q=ms\u2206T"

"s_w=4184J\/kg-K"


"Q=0.01\times4184\times(500-200)=12552J"

work(w)=12552J=12.552KJ

entropy change

"s=\frac{∆Q}{T'}"

"s=\frac{12552}{873}=14.37J/K"